OpenStudy (anonymous):

can you check my answer? what is the remainder when 3x^4+2x^3-x^2+2x-9)/(x+2)?

OpenStudy (anonymous):

i got 5

OpenStudy (solomonzelman):

I am getting 15 as the remainder.

OpenStudy (solomonzelman):

I'll recheck it.

OpenStudy (solomonzelman):

I think it's 15.

OpenStudy (anonymous):

i did it again and your right

OpenStudy (solomonzelman):

Yeah, good work!

OpenStudy (anonymous):

thnks :)

OpenStudy (solomonzelman):

yw

OpenStudy (anonymous):

can u help with one more? @SolomonZelman

OpenStudy (solomonzelman):

I'll try my best.

OpenStudy (anonymous):

ok i hate this question i always get it wrong. what are possible number of positive negative and complex zeros of f(x)=x^6-x^5-x^4+4x^3-12x^2+12?

OpenStudy (solomonzelman):

\[f(x)=x^6-x^5-x^4+4x^3-12x^2+12\] ( just to see it clearly)

OpenStudy (anonymous):

thats right

OpenStudy (anonymous):

positive 3 or 1 negative 2 or 0 complex 4 2 or 0 positive 3 or 1 negative 3 or 1 complex 4 2 or 0 positive 4 2 or 0 negative 2 or 0 complex 6 4 2 or 0 positive 2 or 0 negative 2 or 0 complex 6 4 or 2

OpenStudy (solomonzelman):

What I am think of is this. \[f(x)=x^5(x-1)-x^4+4x^3-12(x^2-1)\]\[f(x)=x^5(x-1)-x^4+4x^3-12(x-1)(x+1)\]\[(x^5+12x+12)(x-1)-x^4+4x^3\]\[(x^5+12x+12)(x-1)-x^3(x+4)\] but factoring would work.

OpenStudy (solomonzelman):

I would suggest plugging in the choices.

OpenStudy (solomonzelman):

I meant but factoring would NOT work....

OpenStudy (anonymous):

im thinking d

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

\[f(x)=x^6-x^5-x^4+4x^3-12x^2+12\], IDK sorry.

OpenStudy (solomonzelman):

If you had one zero given I might have been able to do it.

OpenStudy (anonymous):

what do u think it is??

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (solomonzelman):

Positive 3 doesn't work so not A or B.

OpenStudy (anonymous):

thnk you :)

OpenStudy (solomonzelman):

And positive 4 doesn't work, cuz x^6 is a HUGE number it won't equal zero is x=4

OpenStudy (solomonzelman):

So I therefore think D.

OpenStudy (anonymous):

yay!!!!!!!!! thnks your really the best!!

OpenStudy (solomonzelman):

Yeah, sometimes, I can do the problem, by elimination... my method!!! YW!